If (n+2)/(5) is an integer, what is the remander when (n) is divided by (5)?
Answer by jim_thompson5910(35256)

If (n+2)/(5) is an integer, then n+2 is a multiple of 5, so n+2 = 5k for some integer k

So n+2-2 = n is 2 less than a multiple of 5, which means n = 5k-2

So the remainder is 5-2 = 3